# Can Input Attributions Interpret the Inductive Reasoning Process Elicited in In-Context Learning?

### v1

@@ Abstract
Elucidating the rationale behind neural models’ outputs has been challenging in the machine learning field, which is indeed applicable in this age of large language models (LLMs) and in-context learning (ICL). When it comes to estimating input attributions (IA), ICL poses a new issue of interpreting which example in the prompt, consisting of a set of examples, contributed to identifying the task/rule to be solved. To this end, in this paper, we introduce synthetic diagnostic tasks inspired by the poverty of the stimulus design in inductive reasoning; here, most in-context examples are ambiguous w.r.t. their underlying rule, and one critical example disambiguates the task demonstrated. The question is whether conventional IA methods can identify such an example in interpreting the inductive reasoning process in ICL. Our experiments provide several practical findings; for example, a certain simple IA method works the best, and the larger the model, the generally harder it is to interpret the ICL with gradient-based IA methods. [^1] We will make our data and scripts public upon the publication of this paper.
@@
Can Input Attributions Interpret the Inductive Reasoning Process
Elicited in In-Context Learning?
@@
Figure 1: Overview of our experimental setup. The majority of in-context examples (gray) are ambiguous, supporting either Rule-A of adding two tokens or Rule-B of doubling tokens. A single disambiguating example (blue) reveals that Rule-A is correct. We investigate whether Input Attribution (IA) methods can track such an inductive reasoning process.
@@ 1 Introduction
In past years, input attribution (IA) methods, e.g., gradient norm Simonyan et al. (2014); Li et al. (2016a), have typically been employed in the natural language processing (NLP) field to interpret input–output associations exploited by neural NLP models Vinyals and Le (2015); Li et al. (2016b). Such IA approaches are now somewhat outdated because of the rise of large language models (LLMs) and mechanistic interpretability (MI) communities Olah et al. (2020); Bereska and Gavves (2024), where the general aim is to understand the internal mechanism of LLMs and control their behaviors by intervening in their internal representations or weight parameters rather than elucidating their input-output associations. Despite the enriched scope of modern interpretability research, such rapid progress has missed some intriguing questions bridging the IA and MI eras: in particular, do conventional IA methods still empirically work in the modern NLP setting, specifically in the context of LLMs and in-context learning (ICL)?
In this paper, we perform a systematic, controlled evaluation of IA methods in interpreting the inductive reasoning process of LLMs in ICL Brown et al. (2020)Can IA methods interpret the example in a few-shot prompt that most influenced LLMs’ inductive learning of the task? This question is worth investigating for several reasons. First, input attribution would still serve as a necessary and sufficient explanation in typical practical cases. Some users might simply seek which part of the context is heavily referred to by an LLM system; here, its detailed internal processes identified by MI methods will be rather redundant.
Second, the modern NLP setting, specifically ICL, differs from the conventional tasks where IA methods have been tested. Simply put, in the conventional settings, the task objective is given (e.g., sentiment analysis) in advance, during training, and the IA is employed to just identify the association between input $X$ (i.e., tokens) and output $y$. In contrast, in ICL, the model should (i) first identify the task solely from limited examples demonstrated and then (ii) answer a new question; the former step particularly introduces a new aspect to be interpreted by IA methods — which example, the pair of $(X_{k},y_{k})$ in input, contributed to identifying the task to be solved for a model. This is a type of instance-based interpretation of neural models Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021), and such an in-context, inductive learning process of LLMs attracts interest from a scientific perspective as well Treutlein et al. (2024); Mueller et al. (2024).
@@
  • Certain simpler IA methods (input erasure, gradient norm) frequently outperform other IA methods (e.g., Integrated Gradient) and other interpretability approaches such as observing attention weights, suggesting the particular (dis)advantages of existing IA methods under the ICL×$\times$×LLM settings.
  • Gradient-based and other IA methods did not work stably across different tasks and models, posing their general limitations in interpreting the inductive reasoning part of ICL.
  • Different approaches exhibited different properties against the scaling of the experimental setting; for example, IA methods perform better in many-shot scenarios, whereas a particular baseline (self-answer) tends to outperform IA in larger models.
@@ 2.1 Input attribution (IA) methods
Input attribution (IA) methods Denil et al. (2014); Li et al. (2016a); Poerner et al. (2018); Arras et al. (2019); Yin and Neubig (2022) are one of the most popular and traditional techniques for interpreting and explaining the predictions of machine learning models. Specifically, IA methods determine how much each input feature contributes to a particular prediction; that is, given input tokens $X\coloneqq[x_{1},\ldots,x_{n}]$ and output $y$, the IA methods yield the strength of contribution $S(x_{i})$ of each input $x_{i}$ to the output $y$. Note that the input in ICL consists of several in-context examples (§ 2.2); here, we represent all the tokens in an input sequence as $X$, and the answer to the target question is denoted as $y$. We examine the following four representative IA methods in the ICL×$\times$×LLM context:
@@
IE Li et al. (2016c) measures how erasing a certain token $x_{i}$ from the input prompt makes the target token $y_{t}$ less likely to be generated:
@@
where $\bm{x}_{i}\in\mathbb{R}^{d}$ denotes the $d$-dim input token embedding for the $i$-th token in an input, and $\bm{X}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings. $\bm{X}_{\neg i}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{i-1},\bm{x}_{i+1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings without $\bm{x}_{i}$. We emulate this partial input $\bm{X}_{\neg i}$ by introducing an attention mask to zero-out the attention to $\bm{x}_{i}$ in every layer (thus, the original position information holds). $q(y|\bm{X})$ represents the model’s prediction probability for the token $y_{t}$ given input $\bm{X}$.
@@
GN Simonyan et al. (2014); Li et al. (2016a) calculates the attribution (contribution) score for each input token $x_{i}$ by computing the L1 norm of its gradient of the target token $y_{t}$:
@@
$g(\bm{x}_{i},y_{t};\bm{X})=\nabla_{\bm{x}_{i}}q(y_{t}|\bm{X})\text{.}$ (3)
where $g(\bm{x}_{i},y_{t};\bm{X})\in\mathbb{R}^{d}$ denotes the gradient computed for $\bm{x}_{i}$, given an output $y_{t}$ and input embedding sequence $\bm{X}$.
@@
I×$\times$×G Shrikumar et al. (2017); Denil et al. (2014) takes the dot product of the gradient with the input token embedding $\bm{x}_{i}$:
@@
IG Sundararajan et al. (2017) computes attribution scores by accumulating gradients along a straight path from a baseline input $\bm{X}^{\prime}$ to the actual input $\bm{X}$:
@@
where $\bm{X}^{\prime}\coloneqq[\bm{x}^{\prime}_{1},\ldots,\bm{x}^{\prime}_{n}]$ denotes the sequence of baseline embeddings222 We followed the common practice and employed a sequence of zero vectors as the baseline input. We used an interpretability library captum Kokhlikyan et al. (2020) to calculate the IG score and keep all parameters as default., and $\alpha$ denotes the interpolation coefficient. In practice, the integral is approximated using numerical integration with a finite number of steps.
@@
For the IE, GN, and I×$\times$×G methods, we also adopt the contrastive explanation setting Yin and Neubig (2022). IA methods in this setting measure how much an input token $x_{i}$ influences the model to increase the probability of target token $y_{t}$ while decreasing that of foil token $y_{f}$. A foil token can be defined as an output with an alternative, incorrect generalization (§ 3). Contrastive versions of IE, GN, and I×$\times$×G are defined as follows:
@@
We report the results with contrastive explanation in the main part of this paper § 5.
@@ 2.2 Interpreting in-context learning (ICL)
Here, a model is expected to (i) first induce the underlying function (rule) $f$ from examples $E$ and then generate the final output $f(X_{n+1})$.
Eureka example
[…]
@@
Associative-Recall
a $\mapsto$ 6
g $\mapsto$ 3
w $\mapsto$ 5
g $\mapsto$
3
Key–value pairs are in the prompt.
The task is to output a value associated
with a given key.
@@ B. Generate the distinctive token
A. Multiply the numder of tokens by $n$
@@ B. If the sentence is in title case, then True
Table 1: General format of our inductive reasoning tasks. As a baseline setting, we also set Associative-Recall setting to just memorize key-value mappings. The remaining tasks span from somewhat superficial features to linguistic ones. The disambiguating example (the third one in these examples) determines the correct rule and answer (blue or orange) for the final question from two plausible generalizations shown in the “Potential rules” column.
3 Poverty of the stimulus scenarios as testbeds
We evaluate how well a particular IA method can identify the in-context example $e^{*}$ that most contributes to identifying the task—specifically, whether the IA method assigns the highest attribution score to such an example. However, in real-world tasks, it is generally unclear which in-context example is the most influential in disambiguating the task. Moreovdr, in some cases, the task may be solved without relying on any of the examples (e.g., solved by leveraging prior knowledge). Therefore, these are not suitable as a benchmark to evaluate the interpretability method, and we decided to design a synthetic and controlled task setting.
[…]
@@ 3 Poverty of the stimulus scenarios as testbeds
The few-shot examples are ambiguous as to Rule-A: selecting a character in a particular linear position in an input $X_{i}$; or Rule-B: selecting a character that differs from the others in an input $X_{i}$.
@@
The ambiguity of this task is Rule-A: multiply the numbers of tokens in the input $X_{i}$; or Rule-B: add a certain number of tokens to input $X_{i}$.
@@
This task requires distinguishing whether the controlling condition is the type of verb (Rule A) or the category of the object noun (Rule B). There are two verbs (“like” and “love”) and two categories of the object (city or animal).
@@
The potential rules are Rule-A: whether the main verb in the input $X_{i}$ is in ing-form or not; or Rule-B: whether the first token of input $X_{i}$ is “The” or not.
@@
This task involves two rules: Rule-A: whether there is an adjective in $X_{i}$; or Rule-B: whether $X_{i}$ is presented in title case.
Figure 2: Attribution accuracies for each task/model in the 10-shot setting (thus, the chance rate is 10%; red dotted line). The edit distance and attention baselines are indicated by a black dotted line and gray bar. respectively.
@@ 4.1 Overview
We conduct experiments with different numbers of few-shot examples; specifically, we examined 10-shot, 50-shot, and 100-shot settings to test the robustness of IA methods toward somewhat longer demonstrations.
@@
For each synthetic task, we create 200 different questions with different sets of few-shot examples and a target question. In the LD, AM, VO, TA, and PT tasks, the correct rule is selected out of the two candidates (rules A or B shown in Table 1) in a 1:1 ratio. The position of the most influential (i.e., disambiguating) example $e^{*}$ is controlled to be uniformly distributed by chance. We test IA methods using only the questions that models answered correctly.
Metrics
We report the ratio of questions where the most influential example $e^{*}$ is picked by the highest IA score, i.e., how frequently $\operatorname*{arg\,max}_{e_{i}\in E}S(e_{i})$ is equal to $e^{*}$. Henceforth, we call this score attribution accuracy.
@@
We evaluate five LLMs: Llama-2-7B, Llama-2-13B Touvron et al. (2023), Gemma-2-2B, Gemma-2-9B, and Gemma-2-27B Rivière et al. (2024). As a prerequisite of our experiments, the models should be able to learn the task, i.e., be sufficiently sensitive to the disambiguating example $e^{*}$ and correctly use this to determine the rule. To ensure this ability, we fine-tune these models on each task (see Appendix A), but the conclusions overall did not alter before and after fine-tuning (see Appendix B).

### v2

# Can Input Attributions Interpret the Inductive Reasoning Process in In-Context Learning?

@@ Abstract
Interpreting the internal process of neural models has long been a challenge. This challenge remains relevant in the era of large language models (LLMs) and in-context learning (ICL); for example, ICL poses a new issue of interpreting which example in the few-shot examples contributed to identifying/solving the task. To this end, in this paper, we design synthetic diagnostic tasks of inductive reasoning, inspired by the generalization tests in linguistics; here, most in-context examples are ambiguous w.r.t. their underlying rule, and one critical example disambiguates the task demonstrated. The question is whether conventional input attribution (IA) methods can track such a reasoning process, i.e., identify the influential example, in ICL. Our experiments provide several practical findings; for example, a certain simple IA method works the best, and the larger the model, the generally harder it is to interpret the ICL with gradient-based IA methods.
@@
Can Input Attributions Interpret
the Inductive Reasoning Process in In-Context Learning?
@@
Figure 1: Overview of our experimental setup. The majority of in-context examples (gray) are ambiguous, supporting either Rule A of adding two tokens or Rule B of doubling tokens. A single disambiguating example (blue) reveals that Rule A is correct. We investigate whether input attribution (IA) methods can track such an inductive reasoning process.
@@ 1 Introduction
In past years, input attribution (IA) methods, e.g., gradient norm Simonyan et al. (2014); Li et al. (2016a), have typically been employed in the natural language processing (NLP) field to interpret input–output associations exploited by neural NLP models Vinyals and Le (2015); Li et al. (2016b). Recently, large language models (LLMs) and mechanistic interpretability (MI) research Olah et al. (2020); Bereska and Gavves (2024) have shifted the research focus to understanding the circuits within LLMs by intervening in their internal representations. Despite the enriched scope of research, such rapid progress has missed some intriguing questions bridging the IA and MI eras: in particular, do conventional IA methods still empirically work in the modern NLP setting, specifically in the context of LLMs and in-context learning (ICL)?
In this paper, we revisit IA methods in interpreting LLM-based ICL Brown et al. (2020). Specifically, we assess how well IA methods can track the most influential example in a few-shot examples. This question is worth investigating for several reasons. First, input attribution would still serve as a necessary and sufficient explanation in typical practical cases; some users might simply seek which part of the context is heavily referred to by an LLM system rather than LLMs’ internal processes identified by MI methods. Second, the modern NLP setting, specifically ICL, differs from the conventional settings where IA methods have been tested — identifying the input-output association within a specific test instance ($X_{k}$, $y_{k}$). In contrast, applying IA methods to the entire ICL input (few-shot examples) already entails tracking the learning process as well as the input-output association within a specific target instance. That is, this extended scope includes the interpretation of which example among the demonstrations [$(X_{1},y_{1}),\cdots,(X_{k-1},y_{k-1})$] contributed to identifying the targeted task/rule and then answer a target question $X_{k}$. This is rather a type of instance-based interpretation of neural models Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021), and it has been little explored such interpretation is feasible with IA methods.
@@
  • Gradient norm, the simplest IA method, frequently outperforms other interpretability methods (e.g., integrated gradient), suggesting that the advantage of more recently proposed IA methods does not always generalize in interpreting ICL×$\times$×LLM.
  • Our tested interpretability methods, including simple gradient norm, did not work stably across different tasks and models, posing their general limitations in interpreting ICL with IA methods.
  • Different interpretability methods exhibited different properties with respect to scaling the number of in-context examples or model size; for example, IA methods perform better in many-shot scenarios, whereas a particular baseline interpretability method (i.e., self-answer) works well on larger models.
@@ 2.1 Input attribution (IA) methods
Input attribution (IA) methods are commonly-used techniques for interpreting and explaining the predictions of machine learning models (Denil et al., 2014; Li et al., 2016a; Poerner et al., 2018; Arras et al., 2019, etc.). Specifically, IA methods determine how much each input feature contributes to a particular prediction; that is, given input tokens $X\coloneqq[x_{1},\ldots,x_{n}]$ and output $y$, the IA methods yield the strength of contribution $S(x_{i})$ of each input $x_{i}$ to the output $y$. Note that the input $X$ in ICL consists of several in-context examples (§ 2.2), and the answer to the target question is denoted as $y$. We examine the following four representative IA methods in the ICL×$\times$×LLM context:
@@
IE Li et al. (2016c) measures how impactful erasing a certain token $x_{i}$ from the input prompt is with respect to outputing $y_{t}$:
@@
where $\bm{X}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings, with each $\bm{x}_{i}\in\mathbb{R}^{d}$ being a $d$-dimensional vector corresponding to the $i$-th token in the input. $\bm{X}_{\neg i}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{i-1},\bm{x}_{i+1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings without $\bm{x}_{i}$. We emulate this partial input $\bm{X}_{\neg i}$ by introducing an attention mask to zero-out the attention to $\bm{x}_{i}$ in every layer (thus, the original position information holds). $q(y|\bm{X})$ represents the model’s prediction probability for the token $y_{t}$ given input $\bm{X}$.
@@
GN Simonyan et al. (2014); Li et al. (2016a) calculates the attribution score for each input token $x_{i}$ by computing the L1 norm of its gradient of the target token $y_{t}$:
@@
$g(\bm{x}_{i},y_{t};\bm{X})=\nabla_{\bm{x}_{i}}q(y_{t}|\bm{X})\text{,}$ (3)
where $g(\bm{x}_{i},y_{t};\bm{X})\in\mathbb{R}^{d}$ denotes the gradient of the prediction probability for $y_{t}$ with respect to $\bm{x}_{i}$, under the given input embedding sequence $\bm{X}$.
@@
I×$\times$×G Shrikumar et al. (2017); Denil et al. (2014) takes the dot product of a gradient with the respective token embedding $\bm{x}_{i}$:
@@
IG Sundararajan et al. (2017) is computed by accumulating gradients along a straight path from a baseline input $\bm{X}^{\prime}$ to the actual input $\bm{X}$:
@@
where $\bm{X}^{\prime}\coloneqq[\bm{x}^{\prime}_{1},\ldots,\bm{x}^{\prime}_{n}]$ denotes the sequence of baseline embeddings111 We followed the common practice and employed a sequence of zero vectors as the baseline input. We used an interpretability library captum Kokhlikyan et al. (2020) to calculate the IG score and keep all parameters as default., and $\alpha$ denotes the interpolation coefficient. In practice, the integral is approximated using numerical integration with a finite number of steps.
@@
For the IE, GN, and I×$\times$×G methods, we adopt a contrastive explanation setting, which Yin and Neubig (2022) have shown to be quantitatively superior to the original non-contrastive setting. IA methods in this setting measure how much an input token $x_{i}$ influences the model to increase the probability of target token $y_{t}$ while decreasing that of foil token $y_{f}$. A foil token can be defined as an output with an alternative, incorrect generalization (§ 3). Contrastive versions of IE, GN, and I×$\times$×G are defined as follows:
@@ 2.2 Interpreting in-context learning (ICL)
Here, a model is expected to first induce the underlying function (rule) $f$ from examples $E$ and then generate the final output $f(X_{n+1})$.
Aha example
Interpreting a model’s in-context learning (ICL) involves identifying when, within the input, the model infers the correct rule $f$. To address this aspect, we propose a unique benchmark that features an explicit “aha moment” ($e^{*}$) within the input prompt. At this moment, the correct rule $f$ can be identified by comparing the aha example with one of the other examples in the prompt. Thus, at least, $e^{*}$ should be one of the two most important examples (see evaluation metrics in § 4.2). Note that, to mitigate the potential confusion, we exclude the case of $e^{*}$ being the first example in the demonstration since, in this case, its next example $e_{2}$ can disambiguate the rule and virtually work as the aha example from the perspective of the incremental reasoning process.
@@ B. Generate the distinctive token
A. Add $m$ tokens
     ($m=2$ in this example)
B. Multiply the numder of tokens by $n$
@@ B. If the sentence is in title case, then True
Associative-Recall
a $\mapsto$ 6
g $\mapsto$ 3
w $\mapsto$ 5
g $\mapsto$
3
Key–value pairs are in the prompt.
The task is to output a value associated
with a given key.
Table 1: Formats of our inductive reasoning tasks. As a baseline setting, we also set Associative-Recall setting to just memorize key-value mappings. The remaining tasks span from somewhat superficial features to linguistic ones. The disambiguating example (the third one in these examples) determines the correct rule and answer (blue or orange) for the final question from two plausible generalizations shown in the “Potential rules” column.
3 Problem settings
We evaluate the performance of each IA method in identifying the crucial in-context example $e^{*}$ necessary for defining the task. In real-world tasks, it is generally unclear which in-context example is the most influential in solving the task, and the task may be solved even without relying on any of the examples (e.g., solved by leveraging prior knowledge). Therefore, these are not suitable as a benchmark to evaluate the interpretability method, and we design a synthetic and controlled tasks.
Our setting is the extension of Mueller et al. (2024); we employ a set of ambiguous inductive learning scenarios inspired by the cognitively-motivated LM analyses (McCoy et al. 2020,Warstadt et al. 2020a; inter alia). In these scenarios, a task $f$ is mostly ambiguous in demonstrations $E$ in the sense that several compatible rules exist to explain the transformations $X\mapsto f(X)$. We extend this setting by adding only one disambiguating example $e^{*}$ (“aha example”), which determines the correct rule $f^{*}$ to be unique, and test whether each IA method can identify this special example as long as models correctly employ this clue $e^{*}$ to resolve the problem. For instance, most examples shown in Figure 1 are ambiguous (with gray color) w.r.t. the two possible rules of (i) adding the same token twice or (ii) multiplying the number of tokens by two. This ambiguity is resolved by comparing the aha example $e^{*}$ (blue example in Figure 1) with any one of the other ambiguous examples. As shown in Table 1, we designed the following tasks as a case study:
@@ 3 Problem settings
The few-shot examples are ambiguous as to Rule A: selecting a character in a particular linear position in an input $X_{i}$; or Rule B: selecting a character that differs from the others in an input $X_{i}$.
@@
The ambiguity of this task is Rule A: add a certain number of tokens to input $X_{i}$; or Rule B: multiply the numbers of tokens in the input $X_{i}$.
@@
This task requires distinguishing whether the type of verb (Rule A) or the category of the object noun (Rule B) matters. We employed two verbs (“like” and “love”) and two categories of the object (city or animal).
@@
The potential rules are Rule A: whether the main verb in the input $X_{i}$ is in ing-form or not; or Rule B: whether the first token of input $X_{i}$ is “The” or not.
@@
This task involves two rules: Rule A: whether there is an adjective in $X_{i}$; or Rule B: whether $X_{i}$ is presented in the title case.
Figure 2: Attribution accuracies for each task/model in the 10-shot setting (thus, the chance rate is 20% and 10% for the top-two and top-one metric, respectively; red dotted line). The edit distance and attention baselines are indicated by a black dotted line and gray bar, respectively.
@@ 4.1 Overview
We conducted experiments with different numbers of few-shot examples; specifically, we examined 10-shot, 50-shot, and 100-shot settings to test the robustness of IA methods toward somewhat longer demonstrations.
@@
For each synthetic task, we create 360 different questions with different sets of few-shot examples and a target question. In the LD, AM, VO, TA, and PT tasks, the correct rule is selected out of the two candidates (rules A or B shown in Table 1) in a 1:1 ratio. The position of the most influential (i.e., disambiguating) example $e^{*}$ is assigned according to a uniform distribution over all positions except the first. We test IA methods using only the questions that models answered correctly.
@@
We evaluate five LLMs: Llama-2-7B, Llama-2-13B Touvron et al. (2023), Gemma-2-2B, Gemma-2-9B, and Gemma-2-27B Riviere et al. (2024). As a prerequisite of our experiments, the models should be able to learn the task, i.e., be sufficiently sensitive to the disambiguating example $e^{*}$ and use this to determine the correct rule. To ensure this ability, we fine-tune these models on each task (see Appendix A), but the conclusions overall did not alter before and after fine-tuning (see Appendix B).
4.2 Metrics
We report two accuracy measures: (i) $e^{*}$ is in the top two examples with the highest IA score (top-2 accuracy), and (ii) $e^{*}$ gets the highest IA score among the input examples (top-1 accuracy). Top-2 accuracy is motivated by the fact that models should at least consider the $e^{*}$ plus any other example to identify the correct rule (as described in § 2.2). Top-1 accuracy is motivated from a leave-one-out perspective; excluding the $e^{*}$ significantly hurts the task answerability, while excluding the other examples does not hurt the task ambiguity/complexity.